From: NASA HQ/SpaceRef/NASA Watch
Posted: Wednesday, March 16, 2005
SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2005) to enhance access to related status reports and NASA activities.
All ISS systems continue to function nominally, except those noted previously or below.
After two restart attempts, Elektron was successfully activated on the third. It is currently operating nominally in 50 amp mode. Ground teams will continue to monitor its performance. Additional troubleshooting options, if required, will be addressed following EVA-13 (still scheduled for 3/28). [Today's crew activity schedule was rearranged to accommodate several hours' worth of Elektron-VM troubleshooting to restore functionality of the Russian oxygen generator (there is sufficient O2 onboard the Progress for replenishing if necessary to allow troubleshooting and repair activities with Elektron off). Today's work centered on hydro-resistance checking of the O2 and H2 (hydrogen) lines to locate a suspected obstruction. Also part of the activity was the replacement of the almost depleted nitrogen (N2) supply tank (BPA) with a fresh tank delivered on Progress 17. Work postponed from today includes the planned water transfer from 17P, and potable water sampling tasks. Additional time was also gained by Leroy Chiao's early completion of the Airlock SPCU HX (Service & Performance Checkout Unit Heat Exchanger) and FSS (Fluid Servicer System).]
Early this morning, the external Remote Power Controller switch #17 (RPC-17) tripped open, shutting down CMG-2 (control moment gyroscope #2). Automatic software reconfigured the steering law for the two remaining gyros, CMG-3 & CMG-4. Assessment of the anomaly is underway. [After a similar "Failed Open" trip of RPC-17 on 4/21/2004, CMG-3 & CMG-4 performed nominally for two months, with the Russian ACS (attitude control system) thrusters ready to take over at any moment, until the power switch was replaced on 6/30/04 by Padalka and Fincke on EVA-9B.]
Both crewmembers in turn took their fourth periodic on-orbit hearing assessment (O-OHA) test, a NASA environmental health systems examination to assess the efficacy of acoustic countermeasures. [The O-OHA audiogram test involves minimum audibility measurements for each ear over a wide range of frequencies (0.25-10 kHz) and sound pressure levels, with the crewmembers using individual-specific Prophonics earphones, Bose ANC headsets and the SLM (sound level meter). To conduct the testing, the experimenter is supported by special EarQ software on the MEC (medical equipment computer), featuring an up/down-arrow-operated slider for each test frequency that the crewmember moves to the lowest sound pressure level at which the tone can still be heard. The baseline test is required not later than about Flight Day 14 for each new Expedition and is then performed once per month.]
FE Salizhan Sharipov installed a new 17P-delivered software patch in the Wiener laptop for the BSPN payload server.
CDR Leroy Chiao was scheduled for the CHeCS emergency medical operations OBT (on-board training) drill, a 30-min. exercise to refresh the Crew Medical Officer (CMO)'s acuity in applying ACLS (advanced cardio life support) in an emergency. [Today's computer-based proficiency drill focused on administration of intravenous (IV) fluid infusion and the diagnosis and treatment of a tension pneumothorax. (The latter refers to a collection of gas in the pleural space in the chest resulting in collapse of the lung on the affected side. A tension pneumothorax is a life-threatening condition, where the air within the pleural space is under pressure, displacing mediastinal structures and compromising cardiopulmonary function).]
Sharipov performed the third and last part of the current MBI-8 Profilaktika ("countermeasures") fitness assessment series, first with the usual blood tests (to determine lactate and Creatine Kinase levels in the blood with the AccuSport equipment), then by a physical exercise session on the TVIS treadmill, supported by tagup with a ground specialist. Chiao was available as CMO to assist as required. [The TVIS test is identical to the MO-3 test performed on the treadmill in idling (non-motorized) mode with free choice of speeds within certain specified ranges (idle/walk/slow run/moderate run/fast run/walk/recovery). In addition to the nominal test procedure, MBI-8/Part 3 calls for the use of the TEEM-100M gas analyzer during the test, the blood lactate measurements, and subjective evaluation of physical exertion levels (using the Borg Perceived Exertion Scale, viz., 10 steps from very light over hard and very hard to maximum) during the test. At the end of the Creatine Kinase tests, the results were logged, copied from Cardiocassette-2000 recording to OCA for downlink, and reported to the ground via tagup. The activity was also photo-documented with the Nikon D1 digital camera.]
The crew conducted another standard fit check of the Kazbeks, the contoured shock absorbing seats in the Soyuz 9S descent capsule (SA). [This required them to don their Sokol pressure suits, get in their seats and use a ruler to measure the gap between the top of the head and the top edge of the structure facing the head. The results were reported to TsUP. Kazbek-U couches are designed to withstand g-loads during launch and orbital insertion as well as during reentry and brake-rocket-assisted landing. Each seat has two positions: cocked (armed) and noncocked. In the cocked position, they are raised to allow the shock absorbers to function during touchdown. The fit check assures that the crewmember, whose body gains in length during longer-term stay in zero-G, will still be adequately protected by the seat liners for their touchdown in Kazakhstan. 9S serves as CRV (crew return vehicle) in the event of a contingency.]
Two new items were added to Chiao's "job jar" task list, to be worked on his choosing. [The new tasks are: relocation of an LHA (Light Housing Assembly) in the Node to place two working GLAs (General Luminaire Assemblies) on one power bus for margin of safety in a contingency, and conversion of an old SSC (station support computer) laptop to the new ELC3 (EXPRESS Rack 3 laptop) and the old ELC3 to a spare PCS (portable computer system).]
The FE replaced the fan of the gas analyzer (GA) in the DC1 docking compartment.
At ~11:55am EST, Leroy Chiao conducted a crew/ground debriefing telecon on his successful SPCU HX replacement, supported by a list of questions uplinked overnight.
Salizhan also completed the daily routine maintenance of the SOZh/ECLSS system, along with the regular weekly inspection of the air/liquid condensate separator apparatus (BRPK).
The CDR completed his daily 2.5-hr. physical exercise program on RED (resistive exercise device) CEVIS ergometer bike and TVIS treadmill, while this morning's TVIS exercise for the MBI-8/Part 3 protocol took care of Salizhan's daily workout.
Afterwards, Chiao performed the daily transfer of TVIS and RED exercise data files to the MEC.
Battery 2B2 reconditioning is complete. All six US segment (USOS) battery sets are configured nominally to support channel loads. [A new PPL (pre-positioned load) file will be delivered to update the battery 2B2 SOC (state of charge) calculation, based on data received during the reconditioning. The 2B2 capacity test is planned for after EVA-13.]
As Solar Beta Angle is coming down to less than 10 deg today, ISS attitude will be changed from the current sun-pointing XPOP (x-axis perpendicular to orbit plane) back to earth-fixed LVLH XVV (local vertical/local horizontal x-axis in velocity vector). [The 15-min maneuver on RS (Russian segment) thrusters will take place at 9:07pm EST tonight. Control handover to RS motion control is scheduled for 9:00pm, return to USOS momentum management at 9:40pm.]
Today's CEO (crew earth observations) photo targets, limited in the current XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in ram), were Irrawaddy River Delta, Burma (this overpass provided an opportunity for general oblique photography of this large delta. Looking to the left or right of track for views of the channels and islands in the estuary. Wide-field images of the delta provide useful context for higher resolution data), and Lake Nasser, Toshka Lakes, Egypt (the winter rainy season is ending along the Nile Valley, providing clear weather for images of the Toshka Lakes and Lake Nasser. Monitoring of the increase in area of Toshka Lakes [from both precipitation and Lake Nasser overflow] is important to assess establishment of vegetation and shoreline ecosystems in these new water bodies. Looking to the left of track for the lakes).
CEO photography can be viewed and studied at the websites:
See also the website "Space Station Challenge" at:
To view the latest photos taken by the expedition 10 crew visit:
Expedition 10 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/
Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.
Upcoming Key Events:
ISS Orbit (as of this morning, 6:36am EST [= epoch]):
ISS Altitude History
Apogee height -- Mean Altitude -- Perigee height
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA's Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA's Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.
// end //